Question 4
4.1 An interconnected 50 Hz power system consisting of one area has two
turbine-generator units, rated 600 and 800 MVA, with regulation constants
of 0.04 and 0.06 per unit, respectively, based on their respective ratings.
When each unit carries a 350 MVA steady-state load, let the area load
suddenly increase by 250 MVA. Determine:
4.1.1 The area frequency response characteristic on a 1000 MVA base. (5)
4.1.2 ∆
1
Expert's answer
2021-08-16T02:32:52-0400

"R_{p.u.new}=R_{p.u.old}\dfrac{S_{base.new}}{S_{base.old}}"

"R_1=0.04\dfrac{1000}{600}=0.067"

"R_2=0.06\dfrac{1000}{800}=0.075"

The area frequency response characteristics "\beta" is given by

"\beta=\dfrac{1}{R_1}+\dfrac{1}{R_2}\\
\beta=\dfrac{1}{0.067}+\dfrac{1}{0.075}\\=28.3per\ unit"